OpenStudy (anonymous):

Graph the expression sec^2 x - sec^2 x sin^2 x on your calculator. Determine what constant or single function is equivalent to the given expression.

OpenStudy (anonymous):

You can graph this and if the problem requires you to then you will have to do so, but the answer can be gotten much quicker by realizing that sec^2 times sin^2 is sin^2 over cos^2 which is tan^2. Now you have the identity: sec^2 - tan^2 = 1 So, you will always have "1".

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

This identity comes from sec^2 - (sec^2)(sin^2) = (sec^2)(1 - sin^2) = (sec^2)(cos^2) = (cos^2)/(cos^2) = 1

OpenStudy (anonymous):

One more thing, though.

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

You have to realize that your original expression has a sec in it, so it is undefined at pi/2 and 3pi/2 or pi/2 + kpi. Don't forget the "0" in denominator rule!

OpenStudy (anonymous):

oh okay thank ya!

OpenStudy (anonymous):

You're welcome!